OpenStudy (anonymous):

Let f(x) = x – 2 and g(x) = x^2 – 7x – 9. Find f(g(–1)). 21 3 –3 –21 Please explain, I don't really understand what to do.

OpenStudy (aristo):

first find center function vaule which is g(-1) g(-1)=\[(-1)^{2}-7(-1)-9\] =1+7-9 g(-1)=-1 so the center value is -1 we can now replace g(-1) to -1 the solution now becomes f(-1) f(-1)=-1-2 =-3 the answer is -3